The probability that a football game will go into overtime is 12​%. In 60 randomly selected football​ games, what is the mean and the standard deviation of the number that went into​ overtime? Round to the nearest thousandth.
Answer by stanbon(75887)

The probability that a football game will go into overtime is 12​%. In 60 randomly selected football​ games, what is the mean and the standard deviation of the number that went into​ overtime? Round to the nearest thousandth.
mean = n*p = 60*0.12 = 7.2
std = sqrt(npq) = sqrt(7.2*0.88) = 2.517
-------------------
Cheers,
Stan H.
--------------